Quantrex Quantrex AcademyJEE · NEET · NDA PYQs with solutions Open app
NDA 2026
Exam
NDA
Year
2026
Subject
Mathematics
Chapter
Probability

NDA 2026 · Mathematics · Probability

In an entrance test there are multiple choice questions. There are...

In an entrance test there are multiple choice questions. There are four options for each question, of which only one is correct. The probability that a student knows the answer to a question is 90%90\%. If he gets the correct answer to a question, then what is the probability that he was guessing ?

Options

  1. A3740\dfrac{37}{40}
  2. B3637\dfrac{36}{37}
  3. C137\dfrac{1}{37}
  4. D140\dfrac{1}{40}

Correct answer

C. 137\dfrac{1}{37}

Step-by-step solution

Let E1E_1 be the event that the student knows the answer and E2E_2 be the event that the student guesses the answer.

Let AA be the event that the student gets the correct answer.

From the given information:
P(E1)=90100=910P(E_1) = \dfrac{90}{100} = \dfrac{9}{10}
P(E2)=1−P(E1)=110P(E_2) = 1 - P(E_1) = \dfrac{1}{10}

If the student knows the answer, the probability of getting it correct is 11, so P(A∣E1)=1P(A|E_1) = 1.

If the student guesses, since there are 44 options with only one correct, P(A∣E2)=14P(A|E_2) = \dfrac{1}{4}.

The probability that the student was guessing given that the answer is correct is P(E2∣A)P(E_2|A).

Using Bayes' theorem:
P(E2∣A)=P(E2)P(A∣E2)P(E1)P(A∣E1)+P(E2)P(A∣E2)P(E_2|A) = \dfrac{P(E_2) P(A|E_2)}{P(E_1) P(A|E_1) + P(E_2) P(A|E_2)}

P(E2∣A)=110×14910×1+110×14P(E_2|A) = \dfrac{\dfrac{1}{10} \times \dfrac{1}{4}}{\dfrac{9}{10} \times 1 + \dfrac{1}{10} \times \dfrac{1}{4}}

P(E2∣A)=140910+140=1403740=137P(E_2|A) = \dfrac{\dfrac{1}{40}}{\dfrac{9}{10} + \dfrac{1}{40}} = \dfrac{\dfrac{1}{40}}{\dfrac{37}{40}} = \dfrac{1}{37}

Answer: 137\dfrac{1}{37}
Practice Probability on Quantrex Academy →

More from Probability

Passage: Four dice are rolled. Question: What is the probability that getting a total of… NDA 2026Three events A , B and C are such that A and B are disjoint, A and C are independent, B… NDA 2026Passage: Let A , B , C and D be mutually exclusive and exhaustive events and… NDA 2026Passage: Four dice are rolled. Question: What is the probability that getting a total of… NDA 2026For a Binomial distribution with mean 6 and standard deviation 2 , what is the value of… NDA 2026Passage: A 4-digit number is selected at random formed by using the digits 0 , 1 , 2 , 3… NDA 2026Passage: Let A , B , C and D be mutually exclusive and exhaustive events and… NDA 2026Consider the following statements in respect of the events A , B , C : I. (A B C) ( A B C… NDA 2026Passage: Let A , B , C and D be mutually exclusive and exhaustive events and… NDA 2026The probability that a family owns a laptop is 0.68 ; that it also owns a desktop is 0.56… NDA 2025An urn contains 10 white and 5 red balls. If two balls are drawn at random, then what is… NDA 2025An urn contains 5 white, 6 red and 4 blue balls. Three balls are drawn at random. What is… NDA 2025Under which of the following conditions may binomial distribution be used? I. The number… NDA 2025Three faces of a die are black, two faces are white and one face is red. The die is… NDA 2025A person X speaks the truth 4 out of 5 times and person Y speaks the truth 5 out of 6… NDA 2025An event X can happen with probability p and event Y can happen with probability q .… NDA 2025 Full Probability list Probability JEE Maths PYQsJEE Mathematics PYQs All NDA PYQs